Let X be Binomial(4,0.1). The complement includes zero and one infection: P(X≥2)=1−0.9⁴−4×0.1×0.9³=1−0.6561−0.2916=0.0523. Rounded to three decimals this is 0.052.
More Probability and Statistics questions
GATE 2024 BM Q29– Gaussian noise has mean zero and standard deviation 1 mV. Find P(noise>2 mV or noise<−2…
GATE 2023 BM Q11– Find the difference between the mean and median of the dataset 1, 2, 3, 4, 6, 8.
GATE 2023 BM Q12– For a binomial random variable X, which proposed pair of expectation E(X) and variance…
GATE 2022 BM Q32– In the source’s simplified model, three family members independently have equally likely…
GATE 2022 BM Q50– Six subjects have systolic pressures in mmHg [120,90,100,110,140,130] and corresponding…
GATE 2021 BM Q11– For f X(x)=1qex+e-x to be a valid probability density on the real line, find q.